OpenStudy (anonymous):

How does the graph of f(x) compare to the graph of f(x) =3?

OpenStudy (cwrw238):

do you mean f(x) + 3 or f(x) - 3 ?

OpenStudy (anonymous):

sry +3 How does the graph of f(x) compare to the graph of f(x) + 3?

OpenStudy (anonymous):

Everything on the graph is shifted up by three.

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

Think of it like this: y=f(x) means that any x value is equal to any y value in that function, right?

OpenStudy (anonymous):

yes,,, ok that makes sense!

OpenStudy (anonymous):

so if you add 3 to the function, you will get a y value that has increased by 3. Example: y=f(x) +3. x=1. 1+3=4; therefore y = 4.

OpenStudy (anonymous):

got it! tyyt!!!

OpenStudy (anonymous):

Your welcome!